A refrigerated truck leaves a rest stop traveling at a steady rate of 56 mi/h. A car leaves the same rest stop 1/4 h later following the truck at a steady pace of 64mi/h. How long after the truck leaves the rest stop will the car overtake the truck?
I am wondering how you make an equation for a problem like that and then how you would solve it.

Answer by ankor@dixie-net.com(22740)

A refrigerated truck leaves a rest stop traveling at a steady rate of 56 mi/h.
A car leaves the same rest stop 1/4 h later following the truck at a steady pace of 64mi/hr
How long after the truck leaves the rest stop will the car overtake the truck?
:
Let t = travel time of the truck
then
(t-.25) = travel time of the car
:
When the car overtakes the truck they will have traveled the same distance
Write a distance equation: Dist = speed * time
:
car dist = truck dist
64(t-.25) = 56t
64t - 16 = 56t
64t - 56t = 16
8t = 16
t = 16/8
t = 2 hrs after the truck leaves, the car overtakes him
:
:
Check this by finding the actual distance each traveled; they should be the same.
Car: 64(1.75) = 112 mi
Truck: 56(2) = 112 mi